OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 15x - y = -6 5x - 3y = -13

OpenStudy (anonymous):

5(-15x - 6) - 3y = -13 5(15x + 6) - 3y = -13 5x - 3(15x + 6) = -13 5x - 3(-15x - 6) = -13

OpenStudy (blurbendy):

in the first equation, you need y to equal something so you can substitute it into the second equation. so, isolate y in the top equation

OpenStudy (texaschic101):

-y = -6 - 15x (multiply by -1 to make y positive) y = 15x + 6 now sub 15x + 6 in for y in the 2nd equation and what do you get ?